OpenStudy (anonymous):

help plzzz solve. x^2-14x+58=0 A.{3+7i, 3-7i} B.{2+19i, 2-19i} C.{2+9i, 2-9i} D.{7+3i, 7-3i}

OpenStudy (anonymous):

thank you so much :D

OpenStudy (solomonzelman):

\(\large\color{black}{ x^2-14x+58=0 }\) \(\large\color{black}{ x^2-14x=-58}\) \(\large\color{black}{ x^2-14x+49=-58+49}\) \(\large\color{black}{ x^2-14x+49=-9}\) \(\large\color{black}{ (x-7)^2=-9}\) \(\large\color{black}{ x-7=-±3i}\) \(\large\color{black}{ x=7-±3i}\)

OpenStudy (solomonzelman):

the last thing should be, \(\large\color{black}{ x=7±3i}\)